### v1

# Benchmarking Universal Machine Learning Interatomic Potentials for Supported Nanoparticles: Decoupling Energy Accuracy from Structural Exploration

Abstract
Supported nanoparticle catalysts are widely used in the chemical industry. Computational modeling of supported nanoparticles based on density functional theory (DFT) often involves structural searches of stable local minimum energy configurations and molecular dynamics simulations at finite temperature. These are computationally demanding tasks that are intractable within DFT for large systems. In the last two decades, machine learning interatomic potentials (MLIPs) have been successfully used to substantially increase the size and time scales accessible to simulations that retain DFT accuracy. However, training reliable MLIPs is non-trivial as it requires many costly DFT calculations. Recently, several universal MLIPs (uMLIPs) have been developed, which are trained on large datasets that cover a wide range of molecules and materials. Here, we benchmark the accuracy and the efficiency of these uMLIPs in describing Cu nanoparticles supported on Al2O3 surfaces against our domain-specific DP-UniAlCu model. We find that the MACE-OMAT can reproduce reasonably well the low-energy configurations found in global optimization at an energy accuracy comparable to DP-UniAlCu. Interestingly, the MatterSim-v1.0.0-1M model, which exhibits larger deviations in the binding energies, can find even more stable configurations than the other two models in some supported nanoparticle sizes, showing its capability in structure exploration. For MD simulations, MACE-OMAT and MatterSim-v1.0.0-1M can qualitatively reproduce the mean-squared displacements of Cu atoms (MSDCu) predicted by DP-UniAlCu, albeit at roughly two orders of magnitude higher cost. We demonstrate that the uMLIPs can be very useful in simulating supported nanoparticles even without any fine-tuning, though their reduced efficiency remains a limiting factor for large-scale simulations.
keywords:
Machine Learning Interatomic Potential, Supported Nanoparticle
1 Introduction
Atomic simulation is an indispensable tool in understanding various physical phenomena from an atomic perspective. Heterogeneous catalysis, a complex process that usually involves small molecules and solid surfaces, has benefited from valuable insights obtained from atomic simulation.Chen2020ChemRev The proper description of catalytic reactions and catalyst structures requires a reliable method at the quantum mechanical level. Density functional theory (DFT) has been the workhorse for atomic simulation over the years. However, due to its intensive computational cost, DFT-based simulations cannot satisfy the growing interest in the realistic modeling of catalytic systems that require both extended spatial and time scales.Grajciar2018ChemSocRev
Supported nanoparticle catalyst is one of the most widely used catalysts in the chemical industry.Liu2018ChemRev The computational modeling of supported nanoparticles often involves the structure search of stable configurations at 0 K and the molecular dynamics (MD) simulation at finite temperature. Both the structure search and the MD are computationally demanding and become intractable to DFT when the system size increases. The typical size of a supported nanoparticle synthesized in experiments is in the range of 1-10 nm, where a reasonable structure model of a 2-nm-sized nanoparticle together with the substrate may contain thousands of atoms. The supported nanoparticle structure adopted in DFT-based studies is usually a simplified model that only contains a few dozen atoms. Thus, the computational results are difficult to directly compare with experiments. To understand and design better supported nanoparticle catalysts computationally, an accurate and efficient simulation method is highly desirable.
In the last two decades, machine learning interatomic potentials (MLIPs) have been developed to accelerate atomic simulations, which retain DFT accuracy at a force-field-like cost, and find extensive applications in heterogeneous catalysis.Cheng2024PrecisChem, Omranpour2025ACSCatal, Xia2025ChemSocRev However, the development of an MLIP is a non-trivial task, which requires the user to collect a large amount of ab initio data through many training-validation-application cycles. Also, the MLIP is usually specific to one system, and its transferability is limited. For example, an MLIP trained on one crystalline phase may be useful for exploring configurations of another phase, but the energy and the forces are normally less accurate. Recently, several universal MLIPs (uMLIPs) have emerged, namely, DPA2Zhang2024npjComputMater, DPA3,Zhang2025arXiv MACE-MPBatatia2025JCP, and MatterSimYang2024arXiv_mattersim. These uMLIPs are trained on large datasets that cover a wide range of molecules and materials, such as Materials Project (MP),Jain2013APLMater Open Catalyst Project (OC2M),Tran2023ACSCatal and Open Materials Dataset (OMAT),BarrosoLuque2024arXiv and are expected to apply to a wide range of systems. In general, uMLIPs provide a good foundation to develop an MLIP for a specific system since they can generate reasonable configurations without active learning and can be further fine-tuned to a target accuracy if necessary. MACE-MPBatatia2025JCP has been tested on numerous systems from solid materials to heterogeneous interfaces, while other recent uMLIPs have been mainly tested with solid-state systems. Despite all the tests that have been done, the benchmark of the uMLIPs on supported nanoparticles is still lacking.
In this work, we examined several uMLIPs on simulating supported nanoparticles, where two main application scenarios are demonstrated: (1) global optimization that produces low-energy configurations of nanoparticles serving good initial structures for the further investigation of dynamics, and (2) MD that reveals the dynamics of supported nanoparticles at finite temperature. We choose Cu nanoparticles on Al2O3 surfaces as the benchmark system and set the domain-specific deep potential (DP) model developed in our previous workXu2025ACSNano as the benchmark baseline, which is denoted as DP-UniAlCu in the following sections. The DP-UniAlCu model is trained on a dataset named UniAlCu, including various structures of supported Cu nanoparticles on Al2O3 surfaces. The uMLIPs examined by this work include DPA2 (mptrj and oc20),Zhang2024npjComputMater DPA3 (omat24),Zhang2025arXiv MACE-MP (medium),Batatia2025JCP MACE-OMAT (medium),BarrosoLuque2024arXiv and MatterSim (v1.0.0-1M and v1.0.0-5M).Yang2024arXiv_mattersim
2 Results
Minimum Energy Configurations from Cu1 to Cu21. We commence our examination with geometric optimization-based tasks. Figure 1 shows the binding energy (defined in Equation 1) of Cu1-21 on three Al2O3 surfaces by various MLIPs, which are obtained from global optimization in our previous work.Xu2025ACSNano The three Al2O3 surfaces are modeled by $p(3\times 2)$ $\gamma$-Al2O3(100), $p(2\times 2)$ $\gamma$-Al2O3(110), $p(3\times 2\sqrt{3})$ $\alpha$-Al2O3(0001) periodic slabs (Figure S1). The comparison of the absolute total energy is not accessible as these MLIPs are trained on datasets labelled by different DFT calculation settings. The Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional is used for the reported DFT binding energy results and also the dataset to train our DP-UniAlCu model. MACE and MatterSim-branched uMLIPs are trained on datasets labelled by DFT-PBE using a similar setting as ours (Table S1 and S2), which is the default one named as “MPRelaxSet” used in Materials Project.Jain2013APLMater, Ong2013ComputMaterSci The binding energy results from our DFT setting and “MPRelaxSet” is almost identical with a maximum absolute error of 0.10 eV (Figure S2). Since all MLIPs tested here are trained on DFT-PBE, we are able to examine the accuracy of these MLIPs on the binding energy, where a reliable MLIP should be able to reproduce the binding energy trend across different nanoparticle sizes and surfaces. We noticed the dispersion interaction may affect the absolute values of binding energies of supported nanoparticles when compared to experiments,Wellendorff2015SurfSci, Zhao2024ACSCatal however, the DFT-PBE results presented here only served as a reference for the comparison of different MLIPs and the relative binding strength across different nanoparticle sizes and supporting surfaces is preserved.Xu2025ACSNano Figure 1a-c show the binding energy results of Cu1-21 on $\gamma$-Al2O3(100), $\gamma$-Al2O3(110), and $\alpha$-Al2O3(0001), respectively. The MLIP binding energy results are calculated based on the MLIP total energies of the configurations minimized by the corresponding MLIP model starting from the DFT-minimized configurations. No significant structural difference between DFT and MLIP-minimized configurations is observed. Most models give minimized structures with an averaged atomic displacement smaller than $\sim$0.05 Å compared to the DFT-minimized structures except for DPA2-MPTrj and DPA2-OC2M that give displacements larger than $\sim$0.1 Å (Figure S3). All examined uMLIPs can reproduce the binding energy trend as a function of the nanoparticle size on a single Al2O3 surface, namely, the binding energy on $\gamma$-Al2O3(100) is almost independent of the nanoparticle size while the binding energies on $\gamma$-Al2O3(110) and $\alpha$-Al2O3(0001) decrease significantly with the nanoparticle size. When comparing the binding energies of the same nanoparticle size across different Al2O3 surfaces, some uMLIPs cannot capture the relative binding strength to the supporting surface, where the trend by DFT-PBE from the strongest to the weakest should be $\gamma$-Al2O3(110), $\gamma$-Al2O3(100), and $\alpha$-Al2O3(0001). For example, DPA2-MPTrj predicts a stronger binding on $\alpha$-Al2O3(0001) than on $\gamma$-Al2O3(110), which is inconsistent with DFT-PBE. Figure 1d shows the distribution of binding energy errors through the violin plots, where a good model should have an error distribution centered at zero with a small spread. The binding energy error is defined as the difference between the binding energy predicted by an MLIP and the binding energy by DFT-PBE. MACE-OMAT and MatterSim-v1.0.0-5M exhibit small errors with narrow spreads centered near zero, which reproduce both the relationship between the binding energy and the nanoparticle size and the relative binding strength across three Al2O3 surfaces. It is surprising that MACE-OMAT and MatterSim-v1.0.0-5M are not trained explicitly on any structures of supported nanoparticles but have an accuracy comparable to our DP-UniAlCu model. MACE-MP, which has the same architecture as MACE-OMAT but is trained on a smaller dataset, is less accurate than MACE-OMAT and predicts more negative binding energies on all three Al2O3 surfaces, which can be seen from the average binding energy errors. MatterSim-v1.0.0-1M, which is trained on the same dataset as MatterSim-v1.0.0-5M but has a smaller number of model parameters, leads to a degraded accuracy on binding energy, particularly for $\gamma$-Al2O3(110) with the averaged binding energy error of $\sim$2 eV. The above results indicate that the model performance can be improved either by increasing the dataset size (MACE-MP vs. MACE-OMAT) or by increasing the model size (MatterSim-v1.0.0-1M vs. MatterSim-v1.0.0-5M). To examine this model improvement strategy, we trained a MACE model based on our UniAlCu dataset using the same architecture as MACE-OMAT, which is denoted as the MACE-UniAlCu. The details of the training settings and the model errors on the entire UniAlCu dataset are given in Section S3. MACE-UniAlCu exhibits smaller error spreads for all three Al2O3 surfaces compared to both DP-UniAlCu and MACE-OMAT. The DPA2-MPTrj shows much larger errors compared to other MLIP models. The large errors are also observed for the DPA2-OC2M (Figure S5), which is trained on the OC2M dataset.Chanussot2021ACSCatal The DPA3-OMAT exhibits smaller errors than DPA2-MPTrj and DPA2-OC2M due to its improved architecture. The model architecture can affect the accuracy, as seen from that DPA2-MPTrj is the head trained on the same dataset as MACE-MP, but has larger errors, which is similar in the case of DPA3-OMAT and MACE-OMAT. In addition, the multi-head training strategy used in DPA2 and DPA3, despite utilizing available datasets to the maximum extent, compromises the accuracy of a specific head due to inconsistent DFT settings in the datasets.
Global Optimization of Larger Nanoparticles. In practice, the global minimum configurations of supported nanoparticles are not available and are usually difficult to construct manually. Global optimization can explore low-energy configurations and possibly find the global minimum configurations of supported nanoparticles, where the initial structures are randomly generated, and the final structures can have different geometries but similar energies. A good MLIP should be able to minimize the initial random structures that may have unphysically short interatomic distances and distinguish the low-energy configurations from the high-energy ones. Therefore, global optimization is a good test for the uMLIPs to examine their stability in minimization and their performance in finding low-energy configurations. Based on the performance of uMLIPs on the binding energies of Cu1-21, we examined MACE-OMAT and MatterSim-V1.0.0-1M due to their balance in accuracy and efficiency. At the same time, we tested DP-UniAlCu as a baseline. The global optimization based on the genetic algorithm was performed for several larger nanoparticles, including Cu27, Cu38, Cu47 and Cu55. All the combinations of nanoparticle sizes and Al2O3 surfaces are not included in our UniAlCu dataset, making these global optimization tasks good examinations for the model’s extrapolation ability. The periodic slabs of $p(5\times 3)$ $\gamma$-Al2O3(100), $p(3\times 3)$ $\gamma$-Al2O3(110), $p(5\times 3\sqrt{3})$ $\alpha$-Al2O3(0001) shown in Figure S6 are used to accommodate the nanoparticles. In the global optimization task, the initial structures are randomly generated, and the structures in the following 20 generations are generated based on cut-and-splice and mutation algorithms, where the population size is set to 50. The details of the genetic algorithm are given in Section Methods. Among all obtained MLIP-minimized configurations, 50 structures with the lowest energies are further optimized by DFT. For the three MLIPs examined on one nanoparticle-surface combination, there will be 150 DFT-minimized configurations in total, forming the low-energy ensemble.
Table 1 gives the error distributions on the low-energy ensembles. For nanoparticles on the two $\gamma$-Al2O3 surfaces, our DP-UniAlCu model and the MACE-OMAT model have smaller error spreads compared to MatterSim-v1.0.0-1M. Our DP-UniAlCu is slightly better as the MACE-OMAT model gives more negative binding energies and exhibits larger error spreads, for example, Cu38 on $\gamma$-Al2O3(100). The Mattersim-v1.0.0-1M model gives significantly large errors for Cu27, Cu38, and Cu47 on two $\gamma$-Al2O3 surfaces and predicts decreasing trends in the binding energy, while it gives results of Cu55 closer to DFT. This non-uniform accuracy indicates that the extrapolation ability of a uMLIP can be strongly system-dependent and cannot be guaranteed for consistency across systems, which may be rooted deeply in their training datasets. The DP-UniAlCu and MACE-OMAT models exhibit similar non-uniform accuracies but less severe, for example, they are slightly less accurate for two $\gamma$-Al2O3 surfaces than $\alpha$-Al2O3(0001). Though large offsets are observed, the error spreads by MatterSim-v1.0.0-1M are generally small, indicating the relative stability of the low-energy configurations may pertain. For $\alpha$-Al2O3(0001), all three models exhibit larger error spreads. The error averages by DP-UniAlCu and MACE-OMAT are close to zero, while the error average by MatterSim-v1.0.0-1M is around +2 eV. For Cu47 and Cu55, MACE-OMAT performs better than DP-UniAlCu, showing smaller error spreads.
[…]
Finite-Temperature Dynamics. The dynamic behavior of supported nanoparticles is important for understanding the catalytic activity and stability. Several MD-based tasks are designed to examine uMLIPs. We first tested the mean-squared displacements of Cu atoms (MSDCu, defined in Equation 4) of Cu13 on three Al2O3 surfaces with short-time MD simulations, where ab initio MD (AIMD) trajectories are available for reference. The DP-UniAlCu, MACE-UniAlCu, MACE-OMAT, MatterSim-v1.0.0-1M, MatterSim-v1.0.0-5M, and DPA3-OMAT are tested. For each MLIP, 20 MD simulations starting from the same global minimum configuration but initialized from different velocities by Maxwell-Boltzmann distribution at 800 K are performed for 20 ps with a timestep of 2 fs. For AIMD, only one trajectory is performed with the same settings. The window-averaged MSDCu up to 5 ps are shown in Figure 4a-c. The vertical bars are the standard deviations of 20 MD simulations by MLIPs. The model accuracy can be examined by checking whether the AIMD results fall within the standard deviations of MLIPs. For $\gamma$-Al2O3(100), all models except for MACE-OMAT and DPA3-OMAT give MSDCu results that are close to the AIMD results, which exhibit smaller MSDCu values. For $\gamma$-Al2O3(110), all models give consistent results with AIMD. For $\alpha$-Al2O3(0001), all models except for two MatterSim models produce MSDCu results closer to AIMD, which show larger MSDCu values. The radial distribution functions (RDFs) of Cu-Al and Cu-O pairs are shown in Figure 4d-f and Figure 4g-i, respectively, which are averaged over snapshots every 0.1 ps from 0.2 ps to 20 ps. The RDFs of Cu-Al and Cu-O by DP-UniAlCu, MACE-UniAlCu and MACE-OMAT are in good agreement with the AIMD results, while other models give larger deviations. The RDF of Cu-Cu can be well reproduced by all models (Figure S7). Since the 20-ps-long MD simulations are too short to observe any diffusion of the entire nanoparticle, we performed 1-ns-long MD simulations for Cu13 on three $\alpha$-Al2O3(0001) surfaces with DP-UniAlCu, MACE-UniAlCu, MACE-OMAT, MatterSim-v1.0.0-1M, and DPA3-OMAT to investigate the long-time dynamic behavior. Five independent MD simulations are performed for each MLIP starting from the same global minimum configuration but initialized from different velocities by Maxwell-Boltzmann distribution at 800 K. The results of MSDCu up to 100 ps for Cu13 on three Al2O3 surfaces are shown in Figure 5a-c. The vertical bars are the standard deviations of five MD simulations by MLIPs. Since ns-long MD simulations are not readily accessible to AIMD, the benchmark here only served as a cross-validation across different MLIPs by comparing results against our DP-UniAlCu model. Overall, all MLIPs predict qualitatively consistent mobility trend of Cu atoms on three Al2O3 surfaces, where the MSDCu on $\gamma$-Al2O3(100) and $\alpha$-Al2O3(0001) are comparable and larger than $\gamma$-Al2O3(110). MatterSim-v1.0.0-1M gives slightly larger MSDCu for Cu13 on $\gamma$-Al2O3(100) and $\gamma$-Al2O3(110), and much larger MSDCu for $\alpha$-Al2O3(0001) compared to DP-UniAlCu. MACE-OMAT gives almost identical MSDCu for Cu13 on $\gamma$-Al2O3(110) as DP-UniAlCu while more deviations between them are observed for Cu13 on $\gamma$-Al2O3(100) and $\alpha$-Al2O3(0001), where MACE-OMAT gives smaller MSDCu values. By using the MACE-UniAlCu model, which is trained on the same dataset as DP-UniAlCu, the MSDCu results for $\gamma$-Al2O3(100) and $\alpha$-Al2O2(0001) become closer to those by DP-UniAlCu, however, a smaller MSDCu is observed for $\gamma$-Al2O3(110). DPA3-OMAT exhibits a consistent mobility trend as DP-UniAlCu but gives smaller MSDCu values for all three Al2O3 surfaces. To perform a more quantitative comparison, more simulations with a longer time are necessary, which can be challenging for uMLIPs. The computation performance is shown in Figure 5d, which is taken from the 100-ps simulation of an FCC Cu bulk with 2048 atoms. All models are performed on a single NVIDIA 80G A100 GPU with LAMMPSThompson2022ComputPhysCommun except for MatterSim-v1.0.0-1M with ASELarsen2017JPhysCondensMatter. Our DP-UniAlCu model is about two orders faster than other models, achieving the best balance between accuracy and efficiency here. The reduced efficiency of the uMLIPs arises from their large parameter counts, the overhead associated with message passing, and the substantial cost of tensor products when equivariant features are employed.Batatia2025NatMachIntell Though quantitative results for large size and time scales may not be obtained directly by uMLIPs, they are useful models to sample configuration spaces of interest, which can be distilled by cheaper models realizing long-time simulations.
3 Discussion and Conclusion
In this work, we tested several uMLIPs on supported Cu nanoparticles on Al2O3 surfaces and found that they can be very useful in simulating supported nanoparticles even without any fine-tuning, including the global optimization of low-energy configurations and the MD at finite temperature.
The uMLIPs examined in this work can produce reasonable results on the binding energies of small supported nanoparticles (Cu1-21) reported in our previous study,Xu2025ACSNano where MACE-OMAT shows an accuracy comparable to our domain-specific DP-UniAlCu model. In the global optimization of larger supported nanoparticles (Cu27, Cu38, Cu47, and Cu55), our DP-UniAlCu model behaves better than MACE-OMAT and Mattersim-v1.0.0-1M in finding low-energy minimum configurations. Interestingly, Mattersim-v1.0.0-1M, which shows larger deviations in binding energies, can find even more stable configurations than the other two models in some systems, indicating its exploration ability in the configuration space is not solely characterized by the error metrics on a preset dataset. Alongside the exploration ability of MatterSim-v1.0.0-1M, the non-uniform accuracy across different systems is observed, for instance, the binding energy accuracy in one nanoparticle size is better than another size on the same Al2O3 surface. This non-uniform accuracy may be due to the imbalance in the number of configurations from different systems, where the initial dataset is made up of configurations from one specific system and the configurations of other systems are accumulated in active learning. In terms of global optimization, since the system size of interest is usually accessible to DFT, the MLIP-generated configurations can still be verified by DFT to find the global minimum configurations. The model’s stability in generating diverse configurations in a large region of the configuration space is more important than the accuracy of the energy, as it provides more candidates for DFT verification. A model with better accuracy can improve the performance in the workflow of global optimization since it can produce more low-energy configurations close to the DFT global minimum and reduce the number of DFT calculations, however, the improved accuracy generally comes with an increased computational cost. Though minimization is fast for a single structure and can be made parallel, the computational cost can become intensive for complex systems even by MLIPs, where numerous structures are calculated.
The benchmark on the MD of supported nanoparticles shows that the uMLIPs can produce consistent results of MSDCu and RDF in 20-ps-long simulations compared to AIMD. When it comes to long-time MD simulations, it becomes challenging in two aspects: the lack of DFT reference and the computational cost. The property predicted by MLIPs in the long-time limit can only be cross-validated by different models in the absence of AIMD. MACE-OMAT and MatterSim-v1.0.0-1M exhibit qualitative trends in MSDCu up to 100 ps, which are similar to our DP-UniAlCu model and the MACE-UniAlCu trained on our dataset. To obtain more quantitative results, longer MD simulations are required but not feasible for uMLIPs. Our DP-UniAlCu model is about two orders of magnitude faster than MACE-OMAT and Mattersim-v1.0.0-1M, which makes it a good candidate for long-time MD simulations. Currently, the best tradeoff in accuracy and efficiency is still obtained within the DP architecture. The accuracy of DP can be improved by message passing as demonstrated with DP-MPGao2024PCCP and the efficiency is not degraded as much as switching to uMLIPs. Nevertheless, the uMLIPs can be an alternative method to generate configurations through MD simulations for the initial dataset in the MLIP development instead of the perturbation of a few given structures.Zhang2019PRM For example, distillation is a promising strategy to alleviate the computational performance issue of uMLIPs by training a cheaper model on the configurations sampled and labelled by a uMLIP.
The emergence of uMLIPs provides a good foundation to develop domain-specific MLIPs for complex systems, however, the challenges still exist in reducing the computational cost and integrating fine-tuning and distillation in the current framework of active learning. We expect the benchmarks in this work can provide valuable insights into the applications of uMLIPs for a better understanding of supported nanoparticle catalysts.
4 Methods
The DFT calculations used in binding energy computation and ab initio molecular dynamics are performed with the Vienna Ab initio Simulation Package (VASP).Kresse1996ComputMaterSci, Kresse1996PRB The Perdew-Burke-Ernzerhof (PBE) approximation for the exchange-correlation functionalPerdew1996PRL is used in our UniAlCu dataset and in calculating the binding energies of supported nanoparticles. The MACE-MP, MACE-OMAT, MatterSim-v1.0.0-1M, and MatterSim-v1.0.0-5M are trained on the PBE-based datasets with slightly different parameters. The detailed comparison of input settings can be found in Section S2. The MACE-UniAlCu model is trained on our UniAlCu dataset using the same setting as MACE-OMAT, which is discussed in Section S3.
The simulations by uMLIPs are performed with either LAMMPSThompson2022ComputPhysCommun or ASELarsen2017JPhysCondensMatter through interfaces implemented in the GDPy package.Xu2024GDPy The global optimization is based on the genetic algorithm implemented in the GDPy package,Lee2022JPCC, Xu2024GDPy where the crossover used a cut-and-splice algorithmDeaven1995PRL, Vilhelmsen2012PRL and the mutations include the rattle and mirror operations with equal probabilities. The population size is set to 50 and the number of generations is set to 20. The configurations in the first generation are randomly generated, and those in the following generations are produced by the crossover and mutation methods.
The binding energy $E_{b}$ of a supported nanoparticle is defined as
$E_{b}=E_{{\mathrm{Cu}_{N}}/surf}-E_{surf}-E_{{\mathrm{Cu}_{N}}}$ (1)
where $E_{\mathrm{Cu}_{N}/surf}$ is the potential energy of a Cu nanoparticle on an Al2O3 surface, $E_{surf}$ is the potential energy of a pristine Al2O3 surface, and $E_{\mathrm{Cu}_{N}}$ is the potential energy of the global minimum configuration of CuN in the gas phase.
Two ranking metrics are used to quantify the quality of an MLIP-generated low-energy ensemble. The rank based overlap (RBO)Webber2010ACMTOIS is defined as
$q(S,T,p)=(1-p)\sum_{d=1}^{\infty}p^{d-1}\frac{\lvert S_{1:d}\cap T_{1:d}\rvert}{d},$ (2)
[…]
$\tau=\frac{P-Q}{\sqrt{(P+Q+T)(P+Q+U)}}$ (3)
where $P$ is the number of concordant pairs, $Q$ is the number of discordant pairs, $T$ is the number of ties only in the first list, and $U$ is the number of ties only in the second list.
The mean-squared displacements of Cu atoms (MSDCu) is given by
$MSD(t;r)={<\frac{1}{N}\sum_{i=1}^{N}\lVert r(t+\tau)-r(\tau)\rVert^{2}>}_{\tau},$ (4)
where $N$ is the number of Cu atoms, $r(t)$ is the coordinate at timestep $t$, $\tau$ is the time origin, and $\dotsb>_{\tau}$ denotes the average over all time origins with a lagtime in one trajectory.
{acknowledgement}
The funding for this project was provided by Shell International Exploration and Production Inc., USA. The calculations were performed largely using the Princeton Research Computing resources at Princeton University. The authors also acknowledge the computing resources from the National Energy Research Scientific Computing Center (NERSC) operated under Contract No. DE-AC0205CH11231 using NERSC award ERCAP0021510. The software used in the present work has been developed in the Computational Chemical Science Center “Chemistry in Solution and at Interfaces (CSI)” funded by the USA Department of Energy under award DE-SC0019394. RC was partially supported by this grant, and JX benefited of scientific discussions with the members of CSI.
{suppinfo}
7 figures, 2 tables, notes and references
Figure 1: Several MLIPs are examined for the binding energies of Cu1-21 on three Al2O3 surfaces: (a) $\gamma$-Al2O3(100) (b) $\gamma$-Al2O3(110) (c) $\alpha$-Al2O3(0001) (d) violin plots of errors in binding energy. DP-UniAlCu and MACE-UniAlCu have smaller spreads on the binding energy errors since they are specifically trained on these systems. Among uMLIPs, MACE-OMAT and MatterSim-v1.0.0-5M show similar accuracy as DP-UniAlCu, which reproduce the relative binding strength across three surfaces. Degraded performance is observed for MACE-MP and MatterSim-v1.0.0-1M as the former one is trained on a smaller dataset and the latter one is a model with smaller parameters. DPA2-MPTrj shows large energy differences compared to DFT and DPA3-OMAT has reduced errors due to its improved architecture.
Table 1: Energetics of Global Minimum Configurations (units in eV)
Model DP-UniAlCu MACE-OMAT MatterSim-v1.0.0-1M
Surface Size $avg(\Delta E_{b})\downarrow$ $std(\Delta E_{b})\downarrow$ $avg(\Delta E_{b})\downarrow$ $std(\Delta E_{b})\downarrow$ $avg(\Delta E_{b})\downarrow$ $std(\Delta E_{b})\downarrow$
$\gamma$-Al2O3(100) Cu27 -0.56 0.27 -1.95 0.28 -9.67 0.48
Cu38 -0.99 0.22 -2.59 0.54 -13.74 0.55
Cu47 -1.21 0.32 -2.69 0.25 -16.83 0.40
Cu55 -1.33 0.26 -0.80 0.19 1.48 0.35
$\gamma$-Al2O3(110) Cu27 -0.42 0.30 -1.53 0.38 -8.00 0.42
Cu38 -0.70 0.39 -1.85 0.45 -11.91 0.72
Cu47 -0.96 0.55 -2.00 0.55 -14.74 0.93
Cu55 -1.11 0.45 -0.15 0.16 3.34 0.55
$\alpha$-Al2O3(0001) Cu27 -0.27 0.20 -0.94 0.20 1.72 0.49
Cu38 -0.10 0.23 -0.94 0.22 2.36 0.74
Cu47 0.07 0.47 -1.05 0.23 2.17 0.31
Cu55 -0.51 0.47 -1.61 0.19 2.25 1.08
Table 2: Ranking Metrics on Global Minimum Configurations
Surface Size
[…]
Kendall $\tau\uparrow$
DP-UniAlCu MACE-OMAT MatterSim-v1.0.0-1M
$\gamma$-Al2O3(100) Cu27 q 0.36 0.29 0.02
$\tau$ 0.58 0.26 0.05
Cu38 q 0.41 0.01 0.02
$\tau$ 0.63 -0.42 0.07
Cu47 q 0.70 0.21 0.79
$\tau$ 0.51 0.70 0.46
Cu55 q 0.44 0.30 0.26
$\tau$ 0.48 0.48 0.39
$\gamma$-Al2O3(110) Cu27 q 0.37 0.13 0.06
$\tau$ 0.60 0.51 0.46
Cu38 q 0.75 0.28 0.52
$\tau$ 0.59 0.59 0.56
Cu47 q 0.15 0.02 0.01
$\tau$ 0.21 -0.27 -0.14
Cu55 q 0.46 0.33 0.49
$\tau$ 0.26 0.56 -0.01
$\alpha$-Al2O3(0001) Cu27 q 0.14 0.28 0.01
$\tau$ 0.50 0.61 -0.20
Cu48 q 0.54 0.12 0.05
$\tau$ 0.42 0.12 0.05
Cu47 q 0.43 0.51 0.46
$\tau$ 0.27 0.81 0.61
Cu55 q 0.43 0.25 0.01
$\tau$ -0.06 0.43 -0.21
Figure 2: The low-energy configurations of Cu27 and Cu38 on $\gamma$-Al2O3(100), $\gamma$-Al2O3(110), and $\alpha$-Al2O3(0001) by DFT, DP-UniAlCu (blue), MACE-OMAT (green), and MatterSim-v1.0.0-1M (orange). The DFT configurations are made up of structures obtained from MLIP-based global optimization. Each MLIP provides 50 structures for a given combination of the nanoparticle size and the supporting surface, forming a low-energy ensemble of 150 structures for each system. The arrows indicate the positions of the DFT-minimized global minimum configurations in spectra.
Figure 3: The low-energy configurations of Cu47 and Cu55 on $\gamma$-Al2O3(100), $\gamma$-Al2O3(110), and $\alpha$-Al2O3(0001) by DFT, DP-UniAlCu (blue), MACE-OMAT (green), and MatterSim-v1.0.0-1M (orange). The DFT configurations are made up of structures obtained from MLIP-based global optimization. Each MLIP provides 50 structures for a given combination of the nanoparticle size and the supporting surface, forming a low-energy ensemble of 150 structures for each system. The arrows indicate the positions of the DFT-minimized global minimum configurations in spectra.
Figure 4: For each MLIP, 20 MD simulations starting from the same global minimum configuration of Cu13 but initialized from different velocities by Maxwell-Boltzmann distribution at 800 K are performed for 20 ps with a timestep of 2 fs. For AIMD, only one trajectory is performed with the same settings. (a)-(c) The mean square displacement of Cu atoms. The vertical bars are the standard deviations of 20 MD simulations by MLIPs. (d)-(f) The radial distribution functions of Cu-Al. (g)-(i) The radial distribution functions of Cu-O.
Figure 5: (a)-(c) The mean square displacement of Cu atoms for Cu13 on three Al2O3 surfaces at 800 K. (d) The computational performance of DP-UniAlCu, DPA3-OMAT, MACE-OMAT, MACE-UniAlCu, and MatterSim-v1.0.0-1M models on the simulation of an FCC Cu bulk with 2048 atoms. The “katoms*step/s” is defined as the number of thousands of atoms processed per MD step per second. The inset panels show the initial structures of MD simulations.